A fixed price contract including a profit of 340,000 totals 4.6 million with no reserve. In a progress report the CEV was given as 2.75 million and actual cost was 2.8 million. What is the forecasted (estimated) profit in units of currency using forecasting based on the assumption that cost performance on the contract to date will continue for the remainder of the project?
Answer #1

Given Data:

Total Contract Value of Fixed Price contract (TCV) = $4,600,000

Profit = $340,000

CEV (Contract Earned Value) = $2,750,000

Actual Cost, AC = $28,00,000

Cost Performance Index, CPI = CEV / AC = 2,750,000 / 2,800,000 = 0.982

Project Estimate at Completion, EAC = Total Contract Value / CPI = 46,00,000 / 0.982

= $ 4,684,318

Additional Cost incurred = EAC - TCV = $4,684,318 - $4,600,000 = $84,318 (Over and above the Total Contract value. This will in turn diminish the net profit as shown below)

Estimated Profit = Planned Profit - Additional cost incurred = $340,000 - $84,318 = $255,682

Thus, the estimated profit = $255,682